How do I find 3 ordered pairs and the x and y intercepts of 3x + y = 6?
Teacher's directions:
Provide an x/y chart and show your work in finding 3 ordered pairs. •Find the x and y intercepts of the equation.
This is what I've got done so far.
3x + y = 6 lets say x=0
3 * 0 = 0
y = 6 (intercept at 0,6)
3x + y = 6 lets say y=0
y*0= 0
3*2 = 6
x=2 (intercept at 2,0)
Now, I do not know how to find the third ordered pair.

Answer by oberobic(2304)

Pick and x and calculate the y.
y = -3x + 6
(0,6) is the y-intercept
(2,0) is the x-intercept
.
What do you get if x = 1?
y = -3(1) + 6 = -3+6 = 3, which defines the point (1,3)
.
What do you get if x = -1?
y = -3(-1) + 6 = 3+6 = 9, which defines the point (-1,9)
.
What do you get if x = 3?
y = -3(3) + 6 = -9+6 = -3, which defines the point (3,-3)
.
So you can define an infinite number of points that are on a line.
.
Check the graph to see...

.